Solve:
a) I=P+Prt , for t
b) m=dV(1-e) , for e
c) mu=(m+M)v , for M

Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

Solve:
a) I=P+Prt , for t
t = (I-P)/Pr
--------------------
b) m=dV(1-e) , for e
1-e = m/dV
e = 1-(m/dV)
------------------
c) mu=(m+M)v , for M
m+M = mu/v
M = (mu/v)-m
================
Cheers,
Stan H.

Answer by checkley77(12844)

a) I=P+Prt , for t
Prt=I-P
t=(I-p)/Pr answer.
-----------------------
b) m=dV(1-e) , for e
m=dV-dVe
dVe=dV-m
e=(dV-m)/dV answer.
-----------------------
c) mu=(m+M)v , for M
mu=mv+Mv
Mv=mu-mv
M=m(u-v)/v answer.